find MR , TR ,Q*, P*






1. Q= 25-2p




2. P = 100-10Q




3. 50- 0.5 P




4. 10-0.1Q

1
Expert's answer
2022-03-29T12:12:31-0400

Finding MR for "Q = 25-2P"

"Q= 25-2P"


"TR = QP = (25-2P)P"

"=25P -2P^2"


"MR =\frac{\Delta TR}{P} = 25-4P"


2. P = 100-10Q

"TR = PQ = (100- 10Q)Q"

"= 100Q-10Q^2"

"MR =\frac{\Delta TR}{Q} = 100-20Q"


3. 50- 0.5 P

Finding "P^*" We assume Q = 0

"0 = 50 - 0.5P"

"P^* = \frac{50}{0.5} = 100"


4.10-0.1Q

Finding "Q*" we assume P = 0

"0 = 10-0.1Q"

"P^* = \frac{10}{0.1} = 100"